# Accelerating the Exploration of Thermal Materials via a Synergistic Strategy of DFT and Interpretable Deep Learning

### v9

@@ I Introduction
Kolmogorov-Arnold Networks (KANs) [30] , a novel neural network architecture distinct from Multi-Layer Perceptrons (MLPs), show significant promise in solving ordinary differential equations [31], time series classification [32], and fluid modeling [33]. In some cases, smaller KAN models outperform MLPs [RN165, 30]. Notably, their capabilities in symbolic regression offer interpretability advantages that absent in MLPs [RN165, 30]. As previously mentioned, existing work has successfully used black-box models to accurately predict $\kappa_{\rm L}$. However, achieving accurate $\kappa_{\rm L}$ predictions while simultaneously visualizing the model’s decision-making process, akin to white-box model, remains a significant challenge. The emergence of KANs provides an opportunity to address this issue.
@@
Figure 1 illustrates the framework of our work, which aims to model and predict $\kappa_{\rm L}$ using the white-box deep learning model KAN and compare its performance with that of conventional black-box and white-box ML models. In addition to accuracy, we perform sensitivity analysis to examine interpretability differences across various white-box models and to assess the contributions of individual features to $\kappa_{\rm L}$. This approach further elucidates the relationships between $\kappa_{\rm L}$ and key features, as well as feature-feature correlations. Additionally, Crystal Graph Convolutional Neural Network (CGCNN) [34] is employed to predict critical physical features required for $\kappa_{\rm L}$ modeling with high accuracy, forming a two-stage framework that significantly reduces the time required for screening potential materials. This framework enables rapid qualitative assessments of thermal insulators and conductors. Furthermore, validated through DFT and MD calculations, we successfully identified three thermal conductors and one thermal insulator using this framework. Finally, taking two of these materials as case studies, we leverage DFT to provide in-depth insights into their respective phonon thermal transport mechanisms.
II Results
II.1 Establishment and evaluation of $\kappa_{\rm L}$ models
@@
Model Hyperparameter Range/Value
@@
In our case, KAN was implemented using the pykan-0.0.5 library (github.com/KindXiaoming/pykan), while MLP and XGBoost were built using the PyTorch [38] and scikit-learn [39] libraries, respectively. Initially, a trial-and-error approach was employed to determine approximate ranges for the hyperparameters that might yield good performance for KAN, MLP, and XGBoost. Subsequently, we employed the Optuna [40] library to perform automated hyperparameter optimization for these models. The task of identifying optimal hyperparameters for ML models can be conceptualized as finding the optimal solution to a multivariate optimization problem. It is crucial to acknowledge that hyperparameter optimization often converges to local optima rather than the global optimum, with achieving a global optimum remaining an inherently challenging endeavor. Within the hyperparameter space, multiple local optima may exist, and no algorithm can guarantee that the solution obtained is globally optimal [41]. Nevertheless, it is generally observed that with sufficient iterations, optimization algorithms often produce models with comparable performance across different hyperparameter configurations. This suggests that while achieving global optimality remains a complex challenge, practical implementations typically deliver satisfactory results within reasonable computational effort. The hyperparameter optimization ranges are detailed in Table 1, while the final selected values are presented in Table S1.
@@
Figure 2: Model performance on primary dataset. (a) The blue circular markers represent samples from the training set, while the red square scatters are samples from the test set. (b) KAN model architecture based on the primary dataset.
For regression models, the coefficient of determination ($R^{2}$) provides more information compared to other commonly used criteria [42]. As shown in Figure 2(a), on the primary dataset, XGB slightly outperforms KAN, which in turn outperforms MLP; however, the differences are marginal, and all three models are obviously better than SISSO. Additionally, SISSO exhibits severe errors in a very limited subset of samples with extremely low $\kappa_{\rm L}$ values (approximately $\log(\kappa_{\rm L})<-1$). This phenomenon indicates that SISSO struggles with datasets characterized by class imbalance. SISSO tends to fit more to samples with a higher frequency and a narrower range of target values when dealing with imbalanced data, leading to suboptimal extrapolation performance. As a form of symbolic regression, SISSO inherently faces challenges in addressing sample imbalance. Common approaches to mitigate this issue in symbolic regression include resampling [43] and weighting [43, 44]. However, these methods inevitably impact the model’s interpretability [45, 46]. Therefore, we chose not to compromise SISSO’s interpretability solely for the purpose of enhancing its performance. It is noteworthy that XGBoost outperforms the deep learning models in this experiment, which can be attributed to several factors:
@@
$\quad-0.17\times{10^{7}}\gamma{K_{\rm{i}}}\sqrt[3]{{\Theta{S_{\rm{v}}}}}$ (1)
the linear coefficients in the model are obtained through least-square regression. For the specific physical meanings of the features, please refer to Table 2. Our dataset consists of a total of 10 features, but the SISSO model utilizes only seven of these. Some features, such as the space group number, which are typically not considered to be related to $\kappa_{\rm L}$. SISSO effectively identifies and excludes such extraneous variables. However, it overlooks physical quantities that are theoretically linked to $\kappa_{\rm L}$, such as the thermal expansion coefficient [49, 50], while feature sensitivity analysis indicates that the thermal expansion coefficient is regarded as an important parameter for modeling $\kappa_{\rm L}$ (see Section II.2).
Table 2: Features adopted in the primary dataset.
Definition Notation Unit Range
The space group number for the relaxed structure $SG$ Unitless -
@@
Grüneisen parameter $\gamma$ Unitless [0.00210114, 3.86399]
Heat capacity per cell at constant volume (300K) ${{C}_{\text{v}}}$ $k_{\rm B}$/cell [1.11566, 275.775]
Heat capacity per cell at constant pressure (300K) ${{C}_{\text{p}}}$ $k_{\rm B}$/cell [0.0, 286.301]
Thermal expansion coefficient (300K) $\alpha$ K-1 [$-3.17\times 10^{-6}$, $5.66601\times 10^{-4}$]
[…]
Vibrational free energy per atom (300K) $F$ meV/atom [$-233.268$, 210.145]
Static bulk modulus (300K) ${{K}_{\text{s}}}$ GPa [0.0, 440.158]
@@
In this experiment, the network structure of the KAN model was also determined through automatic hyperparameter optimization via Optuna. The final architecture selected a dimensionality of 7 for the sum operation nodes (i.e., the network structure is $10\times 7\times 1$), with one KAN layer preceding and another following this operation. In the first KAN layer, which connects to the input layer, there are $10\times 7=70$ residual activation functions, while in the second KAN layer, connected to the output layer, the number is $7\times 1=7$. The detailed network architecture is shown in Figure 2(b). The MLP we designed features hidden layers with dimensions of $64\times 64\times 64$. This architecture balances robust pattern recognition capabilities and mitigating issues such as gradient explosion or vanishing gradients. Despite achieving comparable performance, KAN and MLP exhibit significant differences in terms of parameter count, with MLP generally requiring a substantially higher number of parameters. The analytical expressions provided by KAN are not equivalent to the model obtained through data training, unlike SISSO, where the resulting symbolic formula is exactly the model itself. In the KAN Layers, each residual activation function consists of a basis function and a linear combination of several B-spline functions. To symbolize the model, KAN selects the elementary functions with the highest linear correlation to the nodes as substitutes. After symbolization, KAN undergoes further training to determine the affine parameters for each symbolic function. Although this process enhances interpretability, it may slightly reduce performance. Nevertheless, in this experiment, the $R^{2}$ value of the symbolic formula derived from KAN, at 0.9655, remains higher than that obtained by SISSO. The symbolic model constructed by KAN is detailed in Supplementary Information Eq. (1). The analytical expression derived by KAN is much more complex than that of SISSO, as each feature is involved in 7 operations, a consequence of the KAN architecture. Please note that each feature in this formula has undergone Min-Max normalization as part of the feature preprocessing. However, this preprocessing step does not affect the model’s interpretability, as it is easy to revert [51]. Despite the fact that the analytical expressions provided by KAN and SISSO accurately capture the relationship between features and $\kappa_{\rm L}$, there are a few points to consider:
@@
  • Both KAN and SISSO achieve reliable predictive accuracy, but they differ in terms of interpretability. For instance, due to dimensional constraints (see Table S1), SISSO excludes certain features during model construction, and we cannot guarantee that the excluded features are necessarily redundant or useless. Similarly, in KAN, while different features have varying weights, we cannot ensure that all features genuinely influence $\kappa_{\rm L}$. Therefore, it is essential to introduce effective and robust quantitative methods to objectively evaluate the interpretability of these models.
II.2 Deepening physical interpretability through sensitivity analysis

### v10

# Accelerating the Discovery of Materials with Expected Thermal Conductivity via a Synergistic Strategy of DFT and Interpretable Deep Learning

@@ Abstract
Keywords: Lattice Thermal Conductivity; Interpretable Deep Learning; Density Functional Theory; Molecular Dynamics; Sensitivity Analysis.
@@ I Introduction
Kolmogorov-Arnold Networks (KANs) [30] , a novel neural network architecture distinct from Multi-Layer Perceptrons (MLPs), show significant promise in solving ordinary differential equations [31], time series classification [32], and fluid modeling [33]. In some cases, smaller KAN models outperform MLPs [34, 30]. Notably, their capabilities in symbolic regression offer interpretability advantages that absent in MLPs [34, 30]. As previously mentioned, existing work has successfully used black-box models to accurately predict $\kappa_{\rm L}$. However, achieving accurate $\kappa_{\rm L}$ predictions while simultaneously visualizing the model’s decision-making process, akin to white-box model, remains a significant challenge. The emergence of KANs provides an opportunity to address this issue.
@@
Fig. 1 illustrates the framework of our work, which aims to model and predict $\kappa_{\rm L}$ using the white-box deep learning model KAN and compare its performance with that of conventional black-box and white-box ML models. In addition to accuracy, we perform sensitivity analysis to examine interpretability differences across various white-box models and to assess the contributions of individual features to $\kappa_{\rm L}$. This approach further elucidates the relationships between $\kappa_{\rm L}$ and key features, as well as feature-feature correlations. Additionally, Crystal Graph Convolutional Neural Network (CGCNN) [35] is employed to predict critical physical features required for $\kappa_{\rm L}$ modeling with high accuracy, forming a two-stage framework that significantly reduces the time required for screening potential materials. This framework enables rapid qualitative assessments of thermal insulators and conductors. Furthermore, validated through DFT and MD calculations, we successfully identified three thermal conductors and one thermal insulator using this framework. Finally, taking two of these materials as case studies, we leverage DFT to provide in-depth insights into their respective phonon thermal transport mechanisms.
II Methods
II.1 Kolmogorov-Arnold Networks (KANs)
where ${{\phi}_{q,p}}:\left[0,1\right]\to\mathbb{R}$ as well as ${\Phi}_{q}$$\mathbb{R}\to\mathbb{R}$. In this context, $p$ denotes the number of top operators, $q$ denotes the number of bottom operators, and $n$ denotes the number of nodes in the bottom network (which corresponds to the dimensionality of the input features in the input layer), $\mathbf{x}=\left({{x}_{1}},...,{{x}_{n}}\right)$ represents feature vector.
In the original K-A representation theorem, the number of nonlinear layers is limited to 2, and the number of hidden layer nodes is set to $2n+1$, which fixes the network structure to $\left[n,2n+1,1\right]$. However, Liu et al. [30] are not constrained by these limitations. In KANs, both the number of layers and the width of the network are arbitrary, which enhances the feasibility of the K-A representation theorem for ML applications. Therefore, for a KAN with ${n}_{0}$}-D inputs, L layers, and ${{n}_{\text{L}}}=1$-D output, a more precise definition is:
the expression can be simplified to:
where B-spline function is defined as:
$B_{i,0}(x)$ $=\begin{cases}1,&\text{if }t_{i}\leq x<t_{i+1},\\ 0,&\text{otherwise.}\end{cases}$
$B_{i,k}(x)$ $=\frac{x-t_{i}}{t_{i+k}-t_{i}}B_{i,k-1}(x)+\frac{t_{i+k+1}-x}{t_{i+k+1}-t_{i+1}}B_{i+1,k-1}(x).$ (5)
II.2 Sure Independence Screening & Sparsifying Operator (SISSO)
${{\hat{H}}^{\left(\text{m}\right)}}\equiv\left\{I,+,-,\times,\div,\exp,\log,\text{abs},\sqrt{{}}{{,}^{-1}}{{,}^{-2}}{{,}^{-3}}\right\}\left[{{\phi}_{1}},{{\phi}_{2}}\right],$ (6)
${{\Phi}_{i}}=\bigcup\limits_{k}{{{{\hat{h}}}_{k}}^{\left(\text{m}\right)}\left[{{{\hat{\phi}}}_{i}},{{{\hat{\phi}}}_{j}}\right]},\forall{{\hat{h}}_{k}}^{\left(\text{m}\right)}\in{{\hat{H}}^{\left(\text{m}\right)}}\text{ and }\forall{{\phi}_{i}},{{\phi}_{j}}\in{{\Phi}_{i-1}},$ (7)
$\mathbf{\hat{P}}={{\mathbf{d}}_{n}}{{\mathbf{c}}_{n}}={{c}_{0}}+{{c}_{1}}{{d}_{1}}+\cdots+{{c}_{n}}{{d}_{n}},$ (8)
II.3 Sobol indices
${{S}_{i}}=\frac{\text{Var}\left[{{\text{E}}_{\sim i}}\left[Y|{{X}_{i}}\right]\right]}{\text{Var}\left[Y\right]},$ (9)
$S_{i}^{\text{T}}=\frac{{{\text{E}}_{\sim i}}\left[\text{Var}\left[Y|{{\mathbf{X}}_{\sim i}}\right]\right]}{\text{Var}\left[Y\right]}.$ (10)
II.4 Kucherenko indices
Unlike metrics such as Sobol, LIME, and SHAP, the Kucherenko indices accounts for the dependencies among input features. As an extension of the Sobol sensitivity indices, the Kucherenko indices is specifically designed to quantify the sensitivity of model outputs to input variables while considering these interdependencies [41, 40, 42]. The Kucherenko indices essentially involves computing Sobol indices after first employing Copulas [43] (typically Gaussian Copulas) to separate each feature’s marginal distribution from the dependency structure [44], as detailed in Supplementary Table 1. This approach approximates the correlated variables as independent before calculating the Sobol indices. The Gaussian copula is constructed based on the standard normal distribution, and its mathematical expression is:
$C\left({{u_{1}},{u_{2}},\ldots,{u_{d}};\Sigma}\right)={\Phi_{\Sigma}}\left[{{\Phi^{-1}}\left({{u_{1}}}\right),{\Phi^{-1}}\left({{u_{2}}}\right)\ldots,{\Phi^{-1}}\left({{u_{d}}}\right)}\right],$ (11)
In Kucherenko indices, the first-order index $K_{1}$ quantifies the extent to which the variance of the target depends on the variance of a feature in isolation, while the total-effect index $K^{\rm T}$ captures the influence of that feature under interaction effects. A simple example is $y=x_{1}+\cos(x_{2})+\sin(x_{1}x_{2})$, where the sensitivity of the terms $\{x_{1},\cos(x_{2})\}$ with respect to $y$ is reflected by $K_{1}$, whereas the interaction term $\{\sin(x_{1}x_{2})\}$ is captured by $K^{\rm T}$. Notably, $K_{1}>K^{\rm T}$} indicates a strong correlation with other features, while $K^{\rm T}=0$ signifies an exact dependence on other inputs [28, 40].
II.5 Dataset construction & feature preprocessing
$x^{\prime}=\frac{{x-{x_{\min}}}}{{{x_{\max}}-{x_{\min}}}},$ (12)
II.6 Implementation of DFT & GPUMD methods
DFT and AIMD calculations are performed using VASP [52, 53]. The Projector-Augmented Wave (PAW) method [54] and the Perdew-Burke-Ernzerhof (PBE) functional [55] within the Generalized Gradient Approximation (GGA) are employed for electron exchange-correlation. The cutoff energy is set to 500 eV, and the electronic convergence threshold is $10^{-8}$ eV. For the calculation of $\kappa_{\rm L}$, both the phonon BTE method and MD simulations are used. For the stable structures, $\kappa_{\rm L}$ are computed using phono3py [56, 57, 58]. While $\kappa_{\rm L}$ of unstable structures are determined through Graphics Processing Units Molecular Dynamics (GPUMD) [59] AIMD-NVT simulations are carried out in the canonical ensemble (NVT) with a Nosé–Hoover thermostat at 300 K for a duration of 5 ns. A uniformly spaced sampling strategy was adopted, from which 250 frames were selected for training and another 250 frames for testing the NeuroEvolution Potential (NEP) model.
The NEP model employs a feed-forward neural network to establish a mapping between local descriptors and the atomic site energies [60]. In a single hidden layer neural network comprising $N_{\rm neu}$ neurons, the energy $U_{i}$ of atom $i$ is expressed as:
$U_{i}=\sum_{\mu=1}^{N_{\rm neu}}\bm{\omega}_{\mu}^{(1)}\tanh\left(\sum_{\nu=1}^{N_{\rm des}}\bm{\omega}_{\mu v}^{(0)}q_{\nu}^{i}-\mathbf{b}_{\mu}^{(0)}\right)-b^{(1)}.$ (13)
Here, $N_{\rm neu}$ denotes the number of descriptor components, and $q_{\nu}^{i}$ represents the $\nu$-th descriptor component of the $i$-th atom. ${\bm{\omega}}_{\mu\nu}^{\left(0\right)}$ and ${\bm{\omega}}_{\mu}^{\left(1\right)}$ are the weight matrix from the input layer to the hidden layer and the weight vector from the hidden layer to the output node, respectively. $\mathbf{b}_{\mu}^{\left(0\right)}$ and $b^{\left(1\right)}$ correspond to the bias vector in the hidden layer and the bias applied to the node $U_{i}$, respectively. $\tanh{\left(x\right)}$ denotes the nonlinear activation function in the hidden layer [61]. The training of the NEP potential is carried out by optimizing the free parameters through the minimization of a loss function, defined as the weighted sum of the RMSEs of the energies, forces, and virial stresses. The non-equilibrium molecular dynamics (HNEMD) method is employed in GPUMD simulations. In the HNEMD method, an additional driving force ($F_{\rm e}$) is applied to each atom:
$\mathbf{F}_{i}^{\rm ext}=F_{\rm e}\cdot\mathbf{W}_{i},$ (14)
where $F_{\rm e}$ is a driving force parameter with the dimension of inverse length, and $\mathbf{W}_{i}$ is the virial tensor, defined as:
$\mathbf{W}_{i}=\sum_{j\neq i}\mathbf{r}_{ij}\otimes\frac{\partial U_{j}}{\partial\mathbf{r}_{j}i}.$ (15)
$U_{j}$ is the energy of atom $j$ at the given site. The driving force parameter should be systematically tested 3 and a $F_{\rm e}$ value of $1\times 10^{-4}{\AA }^{-1}$ has been tested and found to be optimal for this study. After the driving force was applied to generate the homogeneous heat current, the system was further equilibrated in the NVT ensemble until thermal equilibrium was achieved. The cumulative time average of the instantaneous thermal conductivity, as defined by Eq. (16), was then calculated to obtain $\kappa_{\rm avg}(t)$:
$\kappa_{\rm avg}(t)=\frac{1}{t}\int_{0}^{t}\kappa\left(\tau\right)d\tau.$ (16)
Specific heat and phonon dispersions are calculated using Phonopy [62], while the pCOHP analysis is carried out with the LOBSTER code [63]. Structural visualizations and ELF analyses are performed using VESTA [64]. Furthermore, to complement these results, we present additional validation cases based on Machine Learning Interatomic Potentials (MLIPs) [65] in Supplementary Note 7.
III Results
III.1 Establishment and evaluation of $\kappa_{\rm L}$ models
@@
Figure 2: Architecture of the KAN model based on the full dataset. (a) Input layer; (b) First KAN layer with a total of 70 residual activation functions; (c) Sum-operation nodes layer that connects the two KAN layers and performs summation over the outputs of the first KAN layer; (d) Second KAN layer with 7 residual activation functions; (e) Output scalar corresponding to the predicted value of $\log(\kappa_{\rm L})$.
@@
Model Hyperparameter Domain/Value
@@
In our case, KAN was implemented using the pykan-0.0.5 library (github.com/KindXiaoming/pykan), while MLP and XGBoost were built using the PyTorch [69] and scikit-learn [70] libraries, respectively. Initially, a trial-and-error approach was employed to determine approximate ranges for the hyperparameters that might yield good performance for KAN, MLP, and XGBoost. Subsequently, we employed the Optuna [71] library to perform automated hyperparameter optimization for these models. The task of identifying optimal hyperparameters for ML models can be conceptualized as finding the optimal solution to a multivariate optimization problem. It is crucial to acknowledge that hyperparameter optimization often converges to local optima rather than the global optimum, with achieving a global optimum remaining an inherently challenging endeavor. Within the hyperparameter space, multiple local optima may exist, and no algorithm can guarantee that the solution obtained is globally optimal [72]. Nevertheless, it is generally observed that with sufficient iterations, optimization algorithms often produce models with comparable performance across different hyperparameter configurations. This suggests that while achieving global optimality remains a complex challenge, practical implementations typically deliver satisfactory results within reasonable computational effort. The hyperparameter optimization ranges are detailed in Table 1, while the final selected values are presented in Supplementary Table 2.
Figure 3: Prediction performance of various models upon the full dataset (Table 2). (a) KAN; (b) XGB; (c) MLP; and (d) SISSO. Blue circular markers indicate training samples, while red square markers represent test samples. In each parity plot, the $x$-axis denotes the reference values of $\kappa_{\rm L}$, and the $y$-axis shows the corresponding model predictions.
@@
For regression models, the coefficient of determination ($R^{2}$) provides more information compared to other commonly used criteria [73]. As shown in Fig. 3, on the full dataset, XGB slightly outperforms KAN, which in turn outperforms MLP; however, the differences are marginal, and all three models are obviously better than SISSO. Additionally, SISSO exhibits severe errors in a very limited subset of samples with extremely low $\kappa_{\rm L}$ values (approximately $\log(\kappa_{\rm L})<-1$). This phenomenon indicates that SISSO struggles with datasets characterized by class imbalance. SISSO tends to fit more to samples with a higher frequency and a narrower range of target values when dealing with imbalanced data, leading to suboptimal extrapolation performance. As a form of symbolic regression, SISSO inherently faces challenges in addressing sample imbalance. Common approaches to mitigate this issue in symbolic regression include resampling [74] and weighting [74, 75]. However, these methods inevitably impact the model’s interpretability [76, 77]. Therefore, we chose not to compromise SISSO’s interpretability solely for the purpose of enhancing its performance. It is noteworthy that XGBoost outperforms the deep learning models in this experiment, which can be attributed to several factors:
@@
$-0.17\times{10^{7}}\gamma{K_{\rm{i}}}\sqrt[3]{{\Theta{S_{\rm{v}}}}}.$ (17)
The linear coefficients in Eq. (17) are obtained through least-square regression. For the specific physical meanings of the features, please refer to Table 2. Our dataset consists of a total of 10 features, but the SISSO model utilizes only seven of these. Some features, such as the space group number, which are typically not considered to be related to $\kappa_{\rm L}$. SISSO effectively identifies and excludes such extraneous variables. However, it overlooks physical quantities that are theoretically linked to $\kappa_{\rm L}$, such as the thermal expansion coefficient [80, 81], while feature sensitivity analysis indicates that the thermal expansion coefficient is regarded as an important parameter for modeling $\kappa_{\rm L}$ (see Section III.2).
Table 2: Features adopted in the full dataset.
Definition Notation Unit Domain
The space group number for the relaxed structure $\rm SG$} - -
@@
Grüneisen parameter $\gamma$ - [0.00210114, 3.86399]
Heat capacity per cell at constant volume (300K) ${C}_{\text{v}}$} $k_{\rm B}$ [1.11566, 275.775]
Heat capacity per cell at constant pressure (300K) ${C}_{\text{p}}$} $k_{\rm B}$ [0.0, 286.301]
Thermal expansion coefficient (300K) $\alpha$ K-1 $3.17\times 10^{-6}$, $5.66601\times 10^{-4}$]
Vibrational entropy per atom (300K) ${S}_{\text{v}}$} meV/K [0.0165489, 1.03614]
Vibrational free energy per atom (300K) $F$ meV [233.268-233.268, 210.145]
Static bulk modulus (300K) ${K}_{\rm s}$} GPa [0.0, 440.158]
@@
In this experiment, the network structure of the KAN model was also determined through automatic hyperparameter optimization via Optuna. The final architecture selected a dimensionality of 7 for the sum operation nodes (i.e., the network structure is $10\times 7\times 1$), with one KAN layer preceding and another following this operation. In the first KAN layer, which connects to the input layer, there are $10\times 7=70$ residual activation functions, while in the second KAN layer, connected to the output layer, the number is $7\times 1=7$. The detailed network architecture is shown in Fig. 2. The MLP we designed features hidden layers with dimensions of $64\times 64\times 64$. This architecture balances robust pattern recognition capabilities and mitigating issues such as gradient explosion or vanishing gradients. Despite achieving comparable performance, KAN and MLP exhibit significant differences in terms of parameter count, with MLP generally requiring a substantially higher number of parameters. The analytical expressions provided by KAN are not equivalent to the model obtained through data training, unlike SISSO, where the resulting symbolic formula is exactly the model itself. In the KAN Layers, each residual activation function consists of a basis function and a linear combination of several B-spline functions. To symbolize the model, KAN selects the elementary functions with the highest linear correlation to the nodes as substitutes. After symbolization, KAN undergoes further training to determine the affine parameters for each symbolic function. Although this process enhances interpretability, it may slightly reduce performance. Nevertheless, in this experiment, the $R^{2}$ value of the symbolic formula derived from KAN, at 0.9655, remains higher than that obtained by SISSO. The symbolic model constructed by KAN is detailed in Supplementary Information Eq. (17). The analytical expression derived by KAN is much more complex than that of SISSO, as each feature is involved in 7 operations, a consequence of the KAN architecture. Please note that each feature in this formula has undergone Min-Max normalization as part of the feature preprocessing. However, this preprocessing step does not affect the model’s interpretability, as it is easy to revert [82]. Despite the fact that the analytical expressions provided by KAN and SISSO accurately capture the relationship between features and $\kappa_{\rm L}$, there are a few points to consider:
@@
  • Both KAN and SISSO achieve reliable predictive accuracy, but they differ in terms of interpretability. For instance, due to dimensional constraints (see Supplementary Table 2), SISSO excludes certain features during model construction, and we cannot guarantee that the excluded features are necessarily redundant or useless. Similarly, in KAN, while different features have varying weights, we cannot ensure that all features genuinely influence $\kappa_{\rm L}$. Therefore, it is essential to introduce effective and robust quantitative methods to objectively evaluate the interpretability of these models.
III.2 Deepening physical interpretability through sensitivity analysis